OpenStudy (anonymous):

Find an equation of the line that satisfies the given conditons. Through (1,-6); parallel to the line x+2y=6. How do I find it ?

OpenStudy (anonymous):

find the slope of \(x+2y=6\) first your line will have the same slope

OpenStudy (anonymous):

You have to use the equation y = mx + b m = the slope of the line when you solve x + 2y = 6 you get y = 1/2 x +3 When a line is parallel to that it has the same slope so the slope is 1/2 Now you insert the x,y coordinates that you know 1, -6 You now have -6 = 1/2(1) + b Solve that for that value of b and you have your answer

OpenStudy (anonymous):

okay thank you both so much for your help :)

OpenStudy (anonymous):

well except that it is wrong

OpenStudy (anonymous):

when you solve x + 2y = 6 you get y = 1/2 x +3 is not correct when you solve x + 2y = 6 you get y = -1/2 x +3

OpenStudy (anonymous):

oh okay makes sense since when you subtract the x it turns negative on the other side